### v1

# Using Knowledge Graphs to harvest datasets for efficient CLIP model training

Abstract
Training high-quality CLIP models typically requires enormous datasets, which limits the development of domain-specific models – especially in areas that even the largest CLIP models do not cover well – and drives up training costs. This poses challenges for scientific research that needs fine-grained control over the training procedure of CLIP models. In this work, we show that by employing smart web search strategies enhanced with knowledge graphs, a robust CLIP model can be trained from scratch with considerably less data. Specifically, we demonstrate that an expert foundation model for living organisms can be built using just 10M images. Moreover, we introduce EntityNet, a dataset comprising 33M images paired with 46M text descriptions, which enables the training of a generic CLIP model in significantly reduced time.
1 Introduction
Figure 1: We demonstrate how to harvest datasets for training CLIP models with an improved quality-cost trade-off, either for a generic domain (top) or an expert domain (bottom).
Contrastive Language-Image Pretraining (CLIP) [29] has become a cornerstone for training Vision-Language Models (VLMs). CLIP models learn high-quality visual embeddings and establish a link to the semantic level of brief text descriptions by training on pairs of images and their corresponding text descriptions collected from the web. The features and the link between images and text have been used directly for, e.g., zero-shot classification or text-to-image retrieval, and enable dialogues with visual input, such as in the LLaVA family of models [21, 20, 22]. The link can also be exploited in the opposite direction to enable text-conditional image generation, e.g., Stable Diffusion [28].
Training state-of-the-art CLIP models is computationally expensive. The original CLIP model has seen 12.8B image-text pairs, and later works have scaled this further [10, 8]. This need for scale has limited most of the research to finetuning, which comes with reduced architectural flexibility and control over the data selection. It is particularly problematic for analytic research that demands full control over training to find causes of emergent behavior.
The effort to collect vast datasets is also a key bottleneck for building foundation models for expert domains. Although CLIP models are supposed to be generic and cover most of the world, they are not good enough for use in specific expert domains such as medicine or biology. Building foundation models for expert domains requires an efficient data collection process, taking into account the availability of fewer data samples in these domains.
Our goal is to tackle these challenges from the dataset side while keeping the CLIP algorithm fixed. This strategy is backed by recent literature. For example, Li et al. [18] explored CLIP “along three dimensions: data, architecture, and training strategies” and they stress the “significance of high-quality training data”. For Large Language Models (LLMs), data curation was shown to reduce training time and model size, achieved through heavily filtered publicly available web data and synthetic data [2]. With the dataset creation process, we aim (1) for improved performance in the expert domain of living organisms, in order to demonstrate the creation of expert foundation models; and (2) we aim for a good trade-off between training efficiency and model performance on the broad domain of the visual world, in order to enable compute-efficient from-scratch analysis of fully functional CLIP models.
We built a dataset we named EntityNet, where we leveraged knowledge graphs and targeted web image search. Specifically, from the knowledge graphs Wikidata and WordNet, we collected 135k entities (e.g. eagle) as well as their aliases and descriptions. We extracted entity attributes from Wikidata related to color, partonomy, behavior, and other aspects, which we then used to guide an LLM in generating entity-attribute queries for image search.. For example, from the entity plastic and the attribute small we generated the search term small plastic item.
The resulting EntityNet consists of 33M images paired with 45M alt texts and 613k text labels from the knowledge graphs. The dataset is partitioned into a subset of 10M images of living organisms, capturing high-quality visual and semantic information about the taxonomy of animals, plants, and funghi, as well as a subset of 23M images covering a wide range of categories, such as tools, geographical features, materials, and buildings. Notably, from this process we obtain not only alt texts, but also a link back to the knowledge graph information that was used to create the search query for a given image. We show that this information can be used during training to achieve better performance than by training on alt texts alone. The method of creating our dataset is largely generic and can be applied to other knowledge graphs.
Training on this dataset, we obtain a foundation model that is both specialized on the target expert domain and is also able to understand the overall visual world. In our domain-specific evaluations on iNaturalist and RareSpecies, the model demonstrates robust generalization and clearly surpasses CLIP models trained on much more data (Figure 1). On ImageNet, we demonstrate our dataset to be highly compute efficient and to achieve a performance comparable to models trained 20x longer (Figure 1).
  • We propose a method to automatically create a vision-language dataset based on a given knowledge graph and an image search engine.
  • We apply this method to create the EntityNet dataset, consisting of 33M images paired with 45M alt texts and supplementary text information from the knowledge graphs.
  • We train an expert CLIP model for living organisms on a single 8xL40S machine from scratch in 55 hours. This EntityNet-CLIP is highly specialized in the target expert domain of living organisms, and comparably strong on ImageNet.
  • We evaluate our model and a suite of other CLIP models for object classification, image retrieval, and domain shift robustness. In the expert domain of animals and plants, our model achieves higher performance than models with orders of magnitude more parameters or training data. It is also much stronger than CLIP models that specialize only for this domain. In the generic domain, our model performs remarkably well given the low amount of compute required to train it.
Figure 2: We create a dataset for vision-language pretraining: First, we extract entities from knowledge graphs, then generate attributes and natural types for them. We search for different combinations of entities, attributes, and types in image search engines, and collect alt texts for each image. Finally, we train our model on the combined data.
2 Related work
Datasets. Many recent works investigated ways to build large-scale datasets for multimodal training. Radford et al. [29] train the original CLIP model on a closed set of 400M images, with the model weights being released, but not the data. They build their dataset by collecting image-text pairs, where the text includes frequent terms derived from Wikipedia or WordNet nouns, and approximately balance the classes in the result. As a first approach to create a public dataset of this size, Schuhmann et al. [33] built a dataset with 400 million image-text pairs by filtering HTML data from Common Crawl (LAION-400M) [30]. Their main method of filtering is to remove all image-text pairs that have less than 0.3 similarity estimated by the CLIP model. In a follow-up work, they [34] scaled their approach up one order of magnitude with the multilingual LAION-5B dataset. Xu et al. [48] intended to replicate the original CLIP’s data curation approach. They collected image-text pairs from CommonCrawl and filtered them using Wikipedia and WordNet, then balanced the results. Gadre et al. [10] proposed DataComp, a filtering challenge with a candidate pool of up to 13B image-text pairs from CommonCrawl, where the goal is to filter this candidate pool and to run a fixed training pipeline on the resulting data. They proposed a baseline DataComp-1B dataset with 1.4B pairs filtered with a combination of CLIP score and clustering CLIP embeddings to find images close to ImageNet [7] training examples. Fang et al. [8] trained a Data Filtering Network on an internal dataset of 357M human-verified image-text pairs and finetuned it on a set of public, human-annotated datasets. They filtered 42B candidates into the DFN-5B dataset and trained the current top model of the OpenCLIP leaderboard [14].
These large datasets have mostly replaced smaller datasets like ConceptualCaptions12M (CC12M) [4], which relies on unimodal heuristics and Google Cloud Vision APIs to predict image-text similarity. Another popular small dataset is Yahoo Flickr Creative Commons 15M (YFCC15M), a subset of 15M image-text pairs obtained from the YFCC100M dataset which is based on Flickr [38]. Although many works are mostly concerned with scaling up multimodal datasets and models as much as possible, we aim to improve research on high-quality CLIP models in scenarios where data and compute efficiency is important, for example, for setting up a CLIP model for an expert domain or for scientific analysis of CLIP training.
Stevens et al. [35] aimed for a general vision model named BioCLIP for organismal biology and curated the TreeOfLife-10M dataset based on the Encyclopedia of Life [1], iNaturalist 2021 [41] and BIOSCAN-1M [11]. BioCLIP is trained on a mix of english and latin entity names. For evaluation, they curated the RareSpecies benchmark, which tests generalization to 400 species unseen during training. While the authors of BioCLIP used biological domain knowledge to build their dataset, we rather rely on knowledge graphs and propose a dataset collection method for arbitrary domains.
Training algorithms. Various works are concerned with improving CLIP from the algorithmic side. Li et al. [16] simply train on low resolution first, then finetune on higher resolution later. In another work [17], the authors additionally mask a substantial portion of the image to further reduce the amount of input during training. Zhai et al. [51] propose using a sigmoid loss which reduces the computational load especially in big distributed settings. They follow up [39] by extending the training objective using multiple previously developed techniques, including captioning-based pretraining [45], self-distillation [24] and online data curation [40] into one unified recipe. Vasu et al. [42] enhance their training data with synthetic captions created by an image captioning model and use an ensemble of CLIP teachers to train their model. This way, they can increase the learning efficiency by transferring knowledge from bigger models to their smaller models. Chen et al. [5] evaluate the choice of vision encoder and design a hybrid architecture that improves over vanilla vision transformers when trained under the CLIP framework. Such algorithmic improvements are orthogonal to our research, since they would potentially also improve training on our dataset. In this work, we focus on data improvements and fix the algorithm and architecture choices, since this also allows to easily and fairly compare to a big set of already trained Vision Transformer (ViT) based CLIP models.
Li et al. [18] scale down CLIP and analyse the influence of different data, architecture, and training strategies. They find that especially large models need larger datasets, and data quality plays an important role. They create higher quality datasets by applying CLIP filtering to the 3.4B WebLI dataset [6], while we aim to use a different dataset collection process.
3 Dataset creation
In this section, we outline our dataset creation process, consisting of four main steps: entity extraction, attribute generation, query building, and image search. This process is generally applicable to all visual domains covered by the underlying knowledge graph. In this work, we construct a dataset covering the majority of common visual entities in our world. Within this broad domain, we particularly focus on animals and plants, referred to as the living subset in this section. See Figure 2 for an illustration of the dataset creation process.
3.1 Entity extraction
A high-quality list of visual entities forms the basis for our dataset. We build our list of entities from the Wikidata knowledge graph [43], taking advantage of the hierarchical structure provided by the subclass of relation within Wikidata. For example, the entity dog is a subclass of the pet entity, which in turn is a subclass of domesticated animal entity. Through this hierarchy, we can easily collect all entities under a particular governing or super-entity. First, we manually build a list of 21 super-entities that cover most physical and visual entities in Wikidata. For the living subset, the super-entities are just animal and plant. Examples of non-living super-entities are food, building, or physical tool, with all super-entities listed in the supplementary material. Next, we extract all entities from Wikidata that are linked to at least one of the super-entities through the subclass of relation. For animals and plants, Wikidata also models their biological taxon hierarchy via the parent taxon relation. Because the taxon hierarchy substantially increases the coverage of our living subset, we use it together with the regular subclass hierarchy to extract living entities. Note that we do not want to include named entities (e.g., specific persons). Since Wikidata models these via the instance of relation, our extraction process excludes them by using only the subclass of and parent taxon relations. For every entity, we also download its name, description, aliases, and its number of Wikimedia sitelinks  [^1] The number of Wikimedia sitelinks is a commonly used and high-quality proxy for the popularity of an entity [26]. as additional information. See Table 1 for examples of entities. Finally, we apply two filtering steps: First, we remove all entities with a sitelink count below a predefined threshold, eliminating very rare or low-quality entities which are unlikely to produce strong search results. We find that this filtered list of entities still contains a significant amount of non-visual entities. To address this, we use a LLM to classify each entity as visual or non-visual, discarding the non-visual ones.
For our expert domain, the living subset, we also add all nouns from WordNet [9] that are a subclass of the living thing node, excluding humans, named entities and entities that cannot be seen with the bare eye, e.g., microorganisms.
Finally, we employ heuristic methods to detect and remove potentially offensive entities via a profanity filter.
Table 1: Examples of entities and additional information as extracted from the Wikidata knowledge graph. Name, description, and aliases are used as text labels during training. The number of sitelinks are considered a proxy for an entity’s popularity. The name and aliases are used during search.
Entity Description Sitelinks Aliases
tiger species of big cat 216 tigress, tigers, Panthera tigris
chest box-shaped type of furniture 51 coffer, kist
muscle car type of high-performance car 30 high performance car
3.2 Attribute generation
Besides searching for the entities directly, we also aim to search for variations of them in different contexts, by combining them with attributes. We manually define 6 visual attribute categories we want to search for: Color, Pattern and texture, Parts, Shape and size, Environment, and Other. We extract potential attributes for each entity from the Wikidata knowledge graph and prompt an LLM  [^2] In particular, we use three LLMs and merge their generated attributes: Qwen2.5 7B [49], OpenAI GPT-4o, and OpenAI GPT-4o mini (both accessed via the API at platform.openai.com) with this entity and attribute information to generate a list of visual attributes. For each attribute we also generate a search query combining the attribute itself with the corresponding entity. See Table 2 for examples of generated attributes and search queries. We generate between 1 and 10 attributes per category and generate them for the most popular entities only, as image search engines fail to respect attributes in search queries for rare entities, where they often even struggle to return good results for the entity alone.
Table 2: Examples of attributes and corresponding search queries for different entities as generated by the LLM. Search queries are used for image search and during training.
Entity Category Attribute Search query
rock Pattern and texture porous porous rock
wolf Environment snow wolf in the snow
residence Parts arches arches in residence architecture
garlic Shape and size big big garlic bulb
farm Other tourist tourists visiting a farm
boot Color multicolor multicolored boots
3.3 Query building
For the entities themselves, we use their names and aliases as search queries. We search entity-attribute combinations using the search queries generated by the LLM. We then build another set of queries based on the attributes: First, we determine the natural type for each entity. The natural type of an entity is the super-entity that a human would most likely associate with it, e.g., bird for eagle, or clothing for hat. It is neither too general nor too specific, and can typically be used to disambiguate the entity from other entities with the same name. Here, we use an LLM to select the most fitting super-entity from an entity’s super-class hierarchy as its natural type. We also let the LLM generate a short description as to why the selected natural type is a good fit for the entity and use it during training as a potential text label. We then replace entity mentions in the attribute search queries with their natural types. For example, the attribute query eagle in its nest may turn into bird in its nest, or black BMW M4 into black car.
3.4 Image search and filtering
We execute our search queries using the image search APIs of Bing and Google. Initial tests on the living subset revealed Bing’s search results to be of much higher quality at a lower cost, so we rely solely on the Bing API for all other queries. The image search APIs also provide the URL for the website hosting the image, which we use to collect alt texts from the HTML image tag. After downloading images and alt texts, we perform the following postprocessing steps:
  • Similar to Changpinyo et al. [4], we apply relaxed filtering heuristics. We do not use any multimodal filtering but instead rely on the search engines to provide image-text correspondences. We remove text that is longer than 500 chars or formatted in JSON. We also remove images with an aspect ratio of more than 4 or covering less than 4096 pixels.
  • We deduplicate all downloaded images using the Self-Supervised Descriptor for Image Copy Detection method (SSCD) [27]. For duplicates, we retain the largest image and collect all unique alt texts from the duplicates.
  • We detect duplicates between the images and all evaluation datasets using the same SSCD method.
Our final dataset comprises approximately 33M images and 45M alt texts, obtained from 416k queries. This amounts to 79 images per query and 1.4 alt texts per image on average. The total cost for all image search API calls was around 10,000\$. See Table 3 for an overview over the building blocks and proportions of our dataset.
Table 3: Details of our EntityNet dataset. We show the number of unique elements for each column, e.g. the number of images after deduplication or all unique entity aliases in the respective sets.
Query set Images Queries Entities Aliases Attributes Alt texts Example query
World entity 23M 158k 74k 101k - 23M ship
World entity + attribute 19M 139k 6k 16k 20k 16M small handbag
Living entity 9M 72k 63k 51k - 8M kohlrabi
Living entity + attribute 9M 53k 1k 3k 5k 6M tropical plant
All 33M 416k 135k 149k 23k 45M -
4 Experimental setup
4.1 Training
We trained all models using the standard CLIP loss [29] with a batch size of 8,192 for pretraining and 32,768 during finetuning, along with random resized crop augmentation. We sampled text labels from both the image alt texts and knowledge graph - for each image, 50% of the time we chose a random alt text for a given image, and 50% of the time we chose randomly between search query, aliases, or descriptions of the corresponding entity. Our training code is based on OpenCLIP [14]. Further training details and an example for the text sampling can be found in the supplementary material. We trained all models for 18 epochs using AdamW [23]. Training on 33M images takes $\sim$55 hours on 8 L40s GPUs with 48GB VRAM per GPU.
4.2 Evaluated models
On our EntityNet dataset, we trained ViT CLIP models of size B-32 and B-16 from random initialization. For a comparison with a similarly sized dataset, we also trained models on CC12M by downloading all available URLs, and then detecting and removing duplicates relative to the evaluation datasets using the same procedure as detailed in Section 3.4, obtaining 9.3M images. We finetuned B-32 and B-16 CLIP models trained on DataComp-1B on our dataset to compare finetuning and pretraining performance. In addition to the models trained on EntityNet and CC12M, we evaluate the original OpenAI CLIP [29] and models pretrained on DataComp-M/L/1B, CommonPool-M/L [10], and DFN-5B [8]. We also compare with the domain expert model BioCLIP [35], a ViT-B-16 CLIP model finetuned from OpenAI-CLIP on 10M biological image-text pairs.
4.3 Object classification evaluation
To test the VLMs on object classification we use the same procedure as CLIP [29]. Given an image $I$, class names $C_{1},...,C_{N}$, image encoder $f$ and text encoder $g$, we embed the image using the image encoder $\mathbf{v}=f(I)$. To acquire a text embedding for class $C_{c}$, the CLIP authors started by directly encoding the class names as $\mathbf{w}_{c}=g(C_{c})$, e.g., dog. Alternatively, they created several prompts $P$ using templates, e.g., graffiti of a dog, a photo of the cool dog, etc., then encoded each prompt, and computed the average embedding: $\mathbf{w}_{c}=\sum_{p\in P}g(p)/|P|$. They referred to this approach as using “context prompts”. Finally, given the image and text embeddings, the prediction $p$ is the class which has the highest cosine similarity to the image. We evaluate all models on just encoding the class name, as well as on using the average embedding of the 80 context prompts that the CLIP authors used for ImageNet, and report the higher top-1 accuracy. For zero-shot object classification, we require models to not have trained on the training set of the benchmark, in order to test “generalization to unseen datasets” [29]
Benchmarks in the generic domain. We evaluate on ImageNet [7], a popular image classification benchmark [32]. We use the ILSVRC2012 validation set, which contains 50,000 images from 1,000 classes. The classes include simple objects such as park bench, but also more fine-grained labels like 23 types of terrier dogs, e.g., Staffordshire Bull Terrier. We further evaluate the robustness under distribution shifts on ImageNet-A [13], ImageNet-R [12], ImageNet-Sketch [46], ImageNet-V2 [31], and ObjectNet [3] as proposed by Taori et al. [36]. ImageNet-A contains 7500 samples of 200 ImageNet classes. The samples were adversarially filtered to make ResNet-50s misclassify them, providing a more challenging test. ImageNet-R contains 30000 renditions, such as paintings or embroidery, of 200 ImageNet object classes. ImageNet-Sketch contains 50000 sketches, covering 200 ImageNet classes. ImageNet-V2 replicates the original ImageNet generation process, providing an additional 10000 test images. The object-centered ObjectNet contains 18574 images from 113 ImageNet classes with control over background, rotation and viewpoint.
Benchmarks in the expert domain. We evaluate our models on iNaturalist 2021 [41], a fine-grained species classification benchmark that contains 100k images in the validation set of 10k different species. Similar to Parashar et al. [25], we test models on both the english common name and the latin taxon name. We report the best results over both languages. We further test on the Caltech-UCSD Birds (CUB) [44] dataset, which contains 5,794 images of birds in the original author’s test set, each annotated as one of 200 fine-grained bird species, e.g., grasshopper sparrow. Additionally, we evaluate on the Rare Species benchmark proposed by Stevens et al. [35], that comprises 400 species with 30 images each specifically tailored to assess generalization to unseen taxa. To comply with the benchmark requirements of not seeing the testing 400 species during training, we exclude all entities and queries from our dataset that appear in RareSpecies, using substring matching. As class names, we evaluate all text types proposed by Stevens et al. [35]: combinations of the latin taxonomy and the english common name. Same as in Section 4.3 we evaluate on both the CLIP ImageNet prompt and no prompt, and report the better of both accuracies.
4.4 Retrieval evaluation
We evaluate the COCO Karpathy test split [15], a subset of 5000 samples from the MS-COCO [19] dataset paired with 5 text each. We also evaluate the 1000 samples in the Karpathy test split of Flickr30k [50] annotated with 5 texts per image, as well as on the 3600 image-text pairs in XM3600 [37]. We report the average of image-to-text and text-to-image recall@1 over all datasets.
5 Results
Figure 3: Results on image retrieval and robustness against distribution shifts on ImageNet.
Table 4: Results for training CLIP B-32 and B-16 on our EntityNet dataset from scratch. We mark the best and second best result. To measure training cost, we calculate the total MACs (multiply–accumulate operations) performed during training. We only compare zero-shot results and mark results as “–” if the model does not fulfill the zero-shot requirements.
Arch. Dataset MACs Images in Image- Retrie- Distr. iNat. CUB Rare
(1e18) dataset (M) Net val shift 2021 Species
B-32 CC12M 3.7 9.3 28.6 25.6 18.3 0.7 9.2
B-32 CommonPool-M 2.9 128.0 27.2 20.2 19.8 0.8 10.1
B-32 DataComp-M 2.9 14.0 29.7 19.5 20.5 1.0 16.8
B-32 OpenAI 288.6 400.0 63.4 49.6 48.7 7.4 51.8
B-32 DataComp-1B 295.4 1400.0 69.2 54.0 56.3 12.6 73.8
B-32 EntityNet (ours) 13.1 32.7 61.5 37.2 41.0 26.1 79.5 42.7
B-16 BioCLIP TreeOfLife-10M 61.3 10.4 18.6 0.8 15.4 78.1 38.1
B-16 CommonPool-L 78.2 1280.0 57.8 45.6 47.0 4.1 35.1
B-16 DataComp-L 78.2 140.0 63.1 49.4 51.1 6.1 48.1
B-16 DataComp-1B 791.4 1400.0 73.5 57.4 64.4 15.3 79.0
B-16 OpenAI 784.6 400.0 68.3 52.1 58.6 9.2 56.1
B-16 EntityNet (ours) 36.0 32.7 66.2 39.8 47.4 32.0 85.3 47.1
L-14 OpenAI 3328.4 400.0 75.5 54.3 71.4 12.0 62.9
L-14 DataComp-1B 3338.6 1400.0 79.2 61.8 74.9 21.1 85.5
L-14 DFN-2B 3338.6 2000.0 81.4 64.2 74.8 21.6 86.5
H-14 DFN-5B 22164.0 5000.0 83.4 68.7 76.3 25.1 88.1
Table 5: Results for finetuning the DataComp-1B CLIP model on the EntityNet dataset. We mark the best and second best result. To measure training cost, we calculate the total MACs (multiply–accumulate operations) performed during training.
Arch. Dataset MACs Images in Image Retrie- Distr. iNat CUB
(1e18) dataset (M) -Net val shift 2021
B-32 DataComp-1B (Base model) 295.4 1400.0 69.2 54.0 56.3 12.6 73.8
B-32 EntityNet 13.3 32.7 69.5 50.8 53.3 29.5 83.3
B-32 EntityNet (only living organisms) 4.2 10.2 48.2 31.2 33.3 37.0 87.0
B-16 DataComp-1B (Base model) 791.4 1400.0 73.5 57.4 64.4 15.3 79.0
B-16 EntityNet 36.1 32.7 73.5 52.2 61.0 34.9 86.5
B-16 EntityNet (only living organisms) 11.3 10.2 51.4 34.8 39.2 42.7 90.3
Table 6: Analysis of performance when varying dataset composition, text sampling and dataset size. We mark the best and second best result. To measure training cost, we calculate the total MACs (multiply–accumulate operations) performed during training.
Arch. Dataset MACs Images in Image- Retrie- Distr. iNat. CUB Rare
(1e18) dataset (M) Net val shift 2021 Species
B-32 Everything 13.1 32.7 61.5 37.2 41.0 26.1 79.5 42.7
B-32 No living organisms 9.0 22.5 39.2 32.1 28.0 0.8 6.2 6.9
B-32 Only living organisms 4.1 10.2 36.0 16.5 21.0 28.6 83.2 42.0
B-32 No attribute queries 8.7 21.8 54.8 28.6 33.8 25.6 79.7 39.2
B-32 50% alt text 13.1 32.7 61.5 37.2 41.0 26.1 79.5 42.7
B-32 100% alt text 13.1 32.7 59.1 38.3 38.1 22.9 78.8 39.7
B-32 0% alt text 13.1 32.7 55.7 13.5 35.5 24.2 78.1 29.7
B-32 Full size 13.1 32.7 61.5 37.2 41.0 26.1 79.5 42.7
B-32 1/2 size 6.6 16.4 54.1 30.3 33.6 20.1 74.1 36.6
B-32 1/4 size 3.3 8.2 45.2 23.5 25.7 13.2 64.9 28.0
B-32 1/8 size 1.6 4.1 33.3 16.6 17.7 7.3 47.8 19.4
B-32 1/16 size 0.8 2.0 19.9 9.4 10.0 2.8 27.4 11.8
We evaluate CLIP pretrained from scratch on our EntityNet dataset and CLIP models trained on other datasets. In Figures 1 and 3 we contrast the effectiveness of models with their training cost. We show the results in detail in Table 4. In the generic domain, our models surpass others trained on similarly sized datasets while achieving comparable performance on object classification with models trained 20x longer. On image-text retrieval, our model performs similarly to models trained on the same amount of compute. While our pipeline creates a dataset efficient for understanding objects and their properties, understanding complex scenes still requires learning mainly from the alt texts more than from objects and attributes. In the expert domain, we outperform even the largest CLIP models on the challenging iNaturalist 2021 dataset, which requires classifying images among 10k fine-grained species. Our model also excels on CUB by distinguishing 200 bird species better than all other CLIP models of the same size. Further, when compared to the expert model BioCLIP, trained specifically for organismal biology at a similar training cost, our model demonstrates superior performance. On the Rare Species benchmark, our model classifies the species unseen during training better than the expert biology model, showing the effectiveness of our dataset collection method over a manually designed living organism training set.
We investigate improving existing CLIP models via finetuning in Table 5. The results show that our dataset can be leveraged to create expert CLIP models that outperform both the base model and our model pretrained from scratch on the expert domain. This improvement comes at the cost of trading off some capabilities in the other domains. When finetuning only on the expert domain, we trade off more capabilities, yet obtain even stronger experts.
We further validate and verify our design choices through a component analysis in Table 6. Training separately on the generic and the domain expert part of our dataset reveals that, while the best generic model emerges from training on everything, a slightly better expert model is the result of training only on the expert domain (first table segment). However, generalization to unseen species slightly benefits when training on the full dataset, showing that our generic domain data can enhance generalization capabilities within the expert domain. We also observe that generating and downloading attribute queries contributes to improved performance of the pretrained model.
In the second segment of the table, we evaluate the mixture of alt text and knowledge graph labels used during training. Notably, both training only on alt texts or only on knowledge graph labels mostly performs worse than our 50-50 mix. The exception is image-text retrieval, where training fully on alt text performs slightly better. Potentially, the knowledge graph labels are less useful for learning the matching between longer text queries and images, and more useful for learning fine-grained object classification.
Finally, we reduce the scale of our dataset by powers of two. While the model performance expectedly drops with reduced dataset size, the efficiency of our dataset per datapoint stays high, with the model still reaching 33% accuracy on ImageNet with only 4M images.
6 Limitations
The proposed data harvesting approach assumes that there is a knowledge graph for the target domain and that there is a searchable database with noisy pairing of images and text. Especially, the latter assumption can induce extra effort in some domains. In the medical domain, for example, there are massive amounts of paired image-text data, but the data is not publicly available and not connected to a regular search engine. The search queries must be adapted appropriately to the respective local database. Another limitation is the small, but significant drop in performance on image-text retrieval and classifying ImageNet distribution shifts in the generic domain, when finetuning a large model with EntityNet. First, our dataset by design has a strong focus on the expert domain and trades off some performance in the generic domain during finetuning. Second, our search pipeline finds many clean object-centric images and annotates them with entity information, which tremendously helps understanding object semantics, but to improve efficiency on image-text retrieval in a similar way, one needs to tackle the quality of alt texts and their alignment to the images [47]. Finally, we focused on searching photos, which explains slightly lower accuracy when classifying paintings and sketches – the EntityNet dataset simply contains a lower percentage of such types of images than datasets like CommonPool.
7 Conclusions
We demonstrated how to use knowledge graphs to harvest datasets that are efficient for training CLIP models. Our strategy allows us to create an expert domain dataset with little manual effort, enabling the development of CLIP models that significantly outperform standard models in the expert domain. The expert domain dataset can be used either for training a model from scratch or for finetuning an existing vanilla model. The substantial size and diversity of the expert domain dataset ensures that the good generalization properties of CLIP exist also in the expert domain, in contrast to training with an over-specialized expert dataset.
Furthermore, we demonstrated that the proposed harvesting strategy is also viable to create a common domain dataset, which allows us to achieve a better quality-compute trade-off than training with previous datasets. Future work can use our EntityNet dataset to train CLIP models with all emergent properties much more efficiently, thus allowing for experiments, where training can be controlled. So far, this has been possible only with models of lacking quality.
A Detailed results
We show extensive results on object classification in Table 7. For a more detailed analysis of model capabilities on ImageNet, we split the classes into living (410 classes) and other (590 classes) using WordNet: Since ImageNet labels are built on WordNet nouns, we simply select all labels that are children of the living things node for the living set. On iNaturalist, in addition to the 2021 version, we also evaluate on the 2019 version which contains 3,030 images in the validation set, each annotated with one of 1,010 species. We test with the same protocol as on iNaturalist 2021, testing on both english and latin class names and reporting the best accuracy. We show additional results on retrieval in Table 8 and object classification under distribution shift in Table 9.
Table 7: Detailed object classification results. The table is grouped into training from scratch, finetuning, and analyzing components, the groups are separated by double horizontal lines. Each part is again split by single horizontal lines into groups of same model size or same component analysis. Models marked with $\star$ are finetuned. We mark the best and second best result. To measure training cost, we calculate the total MACs (multiply–accumulate operations) performed during training. We only compare zero-shot results and mark results as “–” if the model does not fulfill the zero-shot requirements.
Arch. Dataset MACs Images in ImageNet iNaturalist CUB Rare
(1e18) dataset 1k Living other 2019 2021 species
[…]
B-32 CC12M 3.7 9.3 28.6 27.6 31.1 2.0 0.7 9.2
B-32 CommonPool-M 2.9 128.0 27.2 20.6 33.5 2.6 0.8 10.1
B-32 DataComp-M 2.9 14.0 29.7 25.5 34.5 3.0 1.0 16.8
B-32 OpenAI 288.6 400.0 63.4 65.5 63.1 10.9 7.4 51.8
B-32 DataComp-1B 295.4 1400.0 69.2 71.2 69.1 16.7 12.6 73.8
B-32 EntityNet 13.1 32.7 61.5 68.5 57.9 38.3 26.1 79.5 42.7
B-16 $\star$ BioCLIP TreeOfLife-10M 61.3 10.4 18.6 44.3 2.6 78.1 38.1
B-16 CommonPool-L 78.2 1280.0 57.8 53.2 62.4 6.9 4.1 35.1
B-16 DataComp-L 78.2 140.0 63.1 61.8 65.3 9.1 6.1 48.1
B-16 DataComp-1B 791.4 1400.0 73.5 75.9 73.2 19.5 15.3 79.0
B-16 OpenAI 784.6 400.0 68.3 71.5 67.4 12.5 9.2 56.1
B-16 EntityNet 36.0 32.7 66.2 73.9 62.0 42.2 32.0 85.3 47.1
L-14 OpenAI 3328.4 400.0 75.5 78.9 74.5 15.2 12.0 62.9
L-14 DataComp-1B 3338.6 1400.0 79.2 82.1 78.4 23.6 21.1 85.5
L-14 DFN-2B 3338.6 2000.0 81.4 83.7 80.9 24.1 21.6 86.5
H-14 DFN-5B 22164.0 5000.0 83.4 85.4 83.2 31.4 25.1 88.1
B-32 DataComp-1B (Base model) 295.4 1400.0 69.2 71.2 69.1 16.7 12.6 73.8
B-32 $\star$ EntityNet 13.3 32.7 69.5 73.6 67.8 41.5 29.5 83.3
B-32 $\star$ EntityNet (only living organisms) 4.2 10.2 48.2 76.3 30.9 49.3 37.0 87.0
B-16 DataComp-1B (Base model) 791.4 1400.0 73.5 75.9 73.2 19.5 15.3 79.0
B-16 $\star$ EntityNet 36.1 32.7 73.5 78.1 71.5 46.7 34.9 86.5
B-16 $\star$ EntityNet (only living organisms) 11.3 10.2 51.4 80.2 33.8 54.3 42.7 90.3
B-32 EntityNet Everything 13.1 32.7 61.5 68.5 57.9 38.3 26.1 79.5 42.7
B-32 EntityNet (no living organisms) 9.0 22.5 39.2 17.5 56.1 1.7 0.8 6.2 6.9
B-32 EntityNet (only living organisms) 4.1 10.2 36.0 68.5 15.4 41.4 28.6 83.2 42.0
B-32 EntityNet (no attribute queries) 8.7 21.8 54.8 63.0 50.4 36.4 25.6 79.7 39.2
B-32 EntityNet 50% alt text 13.1 32.7 61.5 68.5 57.9 38.3 26.1 79.5 42.7
B-32 EntityNet 100% alt text 13.1 32.7 59.1 66.5 55.4 36.4 22.9 78.8 39.7
B-32 EntityNet 0% alt text 13.1 32.7 55.7 64.4 51.5 35.4 24.2 78.1 29.7
B-32 EntityNet Full size 13.1 32.7 61.5 68.5 57.9 38.3 26.1 79.5 42.7
B-32 EntityNet 1/2 size 6.6 16.4 54.1 61.5 50.5 30.7 20.1 74.1 36.6
B-32 EntityNet 1/4 size 3.3 8.2 45.2 52.4 41.8 23.4 13.2 64.9 28.0
B-32 EntityNet 1/8 size 1.6 4.1 33.3 39.6 30.5 14.5 7.3 47.8 19.4
B-32 EntityNet 1/16 size 0.8 2.0 19.9 25.0 18.1 7.6 2.8 27.4 11.8
B-32 EntityNet batch size 2,048 13.1 32.7 59.7 65.4 57.2 31.5 21.0 74.1 38.7
B-32 EntityNet batch size 4,096 13.1 32.7 60.9 67.0 58.0 36.2 24.1 77.6 41.0
B-32 EntityNet batch size 8,192 13.1 32.7 61.5 68.5 57.9 38.3 26.1 79.5 42.7
B-32 EntityNet batch size 16,384 13.2 32.7 60.5 68.2 56.6 38.8 26.9 81.2 42.1
B-32 EntityNet batch size 32,768 13.3 32.7 58.6 67.0 54.2 39.5 26.3 81.3 41.7
Table 8: Detailed retrieval results. The table is grouped into training from scratch, finetuning, and analyzing components, the groups are separated by double horizontal lines. Each part is again split by single horizontal lines into groups of same model size or same component analysis. Models marked with $\star$ are finetuned. We mark the best and second best result. To measure training cost, we calculate the total MACs (multiply–accumulate operations) performed during training.
Arch. Dataset MACs Images in Retrieval COCO test F30K test XM3600 test
(1e18) dataset Average I2T T2I I2T T2I I2T T2I
B-32 CC12M 3.7 9.3 25.6 22.4 15.2 37.2 27.1 26.1 25.5
B-32 CommonPool-M 2.9 128.0 20.2 18.3 11.2 29.9 18.9 23.6 19.6
B-32 DataComp-M 2.9 14.0 19.5 17.1 11.0 26.0 18.0 23.6 21.5
B-32 OpenAI 288.6 400.0 49.6 50.1 30.5 77.5 58.8 43.4 37.2
B-32 DataComp-1B 295.4 1400.0 54.0 53.5 37.1 78.8 61.1 48.3 45.3
B-32 EntityNet 13.1 32.7 37.2 32.8 22.2 52.3 37.8 40.6 37.3
B-16 $\star$ BioCLIP TreeOfLife-10M 61.3 10.4 0.8 0.4 0.2 0.9 0.6 1.8 1.1
B-16 CommonPool-L 78.2 1280.0 45.6 44.4 28.8 68.3 51.0 42.1 39.2
B-16 DataComp-L 78.2 140.0 49.4 48.7 32.2 73.5 55.1 44.7 42.1
B-16 DataComp-1B 791.4 1400.0 57.4 57.5 40.2 84.9 67.3 47.9 46.5
B-16 OpenAI 784.6 400.0 52.1 52.5 33.1 81.9 62.0 43.8 39.3
B-16 EntityNet 36.0 32.7 39.8 36.0 25.2 57.1 43.3 39.9 37.1
L-14 OpenAI 3328.4 400.0 54.3 56.3 36.5 85.1 65.2 44.5 38.4
L-14 DataComp-1B 3338.6 1400.0 61.8 63.2 45.8 89.5 73.4 50.4 48.6
L-14 DFN-2B 3338.6 2000.0 64.2 65.7 48.6 89.6 75.3 53.6 52.7
H-14 DFN-5B 22164.0 5000.0 68.7 72.3 53.9 93.0 80.2 57.6 55.4
B-32 DataComp-1B (Base model) 295.4 1400.0 54.0 53.5 37.1 78.8 61.1 48.3 45.3
B-32 $\star$ EntityNet 13.3 32.7 50.8 48.1 34.4 72.1 57.1 47.8 45.6
B-32 $\star$ EntityNet (only living organisms) 4.2 10.2 31.2 28.0 19.7 44.8 33.7 30.9 29.9
B-16 DataComp-1B (Base model) 791.4 1400.0 57.4 57.5 40.2 84.9 67.3 47.9 46.5
B-16 $\star$ EntityNet 36.1 32.7 52.2 50.4 36.4 75.6 59.4 47.3 44.3
B-16 $\star$ EntityNet (only living organisms) 11.3 10.2 34.8 31.8 22.8 51.7 37.1 33.5 32.0
B-32 EntityNet Everything 13.1 32.7 37.2 32.8 22.2 52.3 37.8 40.6 37.3
B-32 EntityNet (no living organisms) 9.0 22.5 32.1 28.6 18.7 46.3 33.6 33.2 32.3
B-32 EntityNet (only living organisms) 4.1 10.2 16.5 12.8 9.8 23.4 15.3 19.0 18.6
B-32 EntityNet (no attribute queries) 8.7 21.8 28.6 23.6 16.1 41.3 27.3 32.3 30.8
B-32 EntityNet 50% alt text 13.1 32.7 37.2 32.8 22.2 52.3 37.8 40.6 37.3
B-32 EntityNet 100% alt text 13.1 32.7 38.3 35.2 23.2 53.4 39.8 40.5 38.0
B-32 EntityNet 0% alt text 13.1 32.7 13.5 8.7 6.1 19.2 11.7 18.3 17.0
B-32 EntityNet Full size 13.1 32.7 37.2 32.8 22.2 52.3 37.8 40.6 37.3
B-32 EntityNet 1/2 size 6.6 16.4 30.3 27.0 17.6 42.7 29.7 33.3 31.6
B-32 EntityNet 1/4 size 3.3 8.2 23.5 19.5 13.0 31.1 22.9 27.6 27.2
B-32 EntityNet 1/8 size 1.6 4.1 16.6 12.8 8.7 20.7 14.2 21.9 20.9
B-32 EntityNet 1/16 size 0.8 2.0 9.4 7.1 5.0 10.0 7.1 13.9 13.5
B-32 EntityNet batch size 2,048 13.1 32.7 35.6 30.8 20.9 50.2 37.2 38.2 36.4
B-32 EntityNet batch size 4,096 13.1 32.7 36.4 31.6 22.1 51.7 37.0 39.1 37.2
B-32 EntityNet batch size 8,192 13.1 32.7 37.2 32.8 22.2 52.3 37.8 40.6 37.3
B-32 EntityNet batch size 16,384 13.2 32.7 35.8 32.2 21.8 50.5 36.9 37.2 36.1
B-32 EntityNet batch size 32,768 13.3 32.7 34.5 31.0 20.8 48.9 34.5 36.2 35.3
Table 9: Detailed results on ImageNet distribution shifts. The table is grouped into training from scratch, finetuning, and analyzing components, the groups are separated by double horizontal lines. Each part is again split by single horizontal lines into groups of same model size or same component analysis. Models marked with $\star$ are finetuned. We mark the best and second best result. To measure training cost, we calculate the total MACs (multiply–accumulate operations) performed during training. INet: ImageNet.
Arch. Dataset MACs Images in INet Ave- INet INet Object- INet INet
(1e18) dataset (M) 1K rage V2 R Net Sketch A
[…]
B-32 CC12M 3.7 9.3 28.6 18.3 24.2 34.5 12.1 16.0 4.7
B-32 CommonPool-M 2.9 128.0 27.2 19.8 22.5 33.0 20.9 18.4 4.3
B-32 DataComp-M 2.9 14.0 29.7 20.5 24.4 34.0 19.7 19.3 4.9
B-32 OpenAI 288.6 400.0 63.4 48.7 56.0 69.4 44.2 42.3 31.5
B-32 DataComp-1B 295.4 1400.0 69.2 56.3 60.8 78.2 55.2 56.8 30.5
B-32 EntityNet 13.1 32.7 61.5 41.0 53.6 58.8 32.6 45.0 14.9
B-16 $\star$ BioCLIP TreeOfLife-10M 61.3 10.4 18.6 15.4 17.7 16.0 3.2 7.3 32.9
B-16 CommonPool-L 78.2 1280.0 57.8 47.0 50.0 68.4 49.1 45.9 21.7
B-16 DataComp-L 78.2 140.0 63.1 51.1 55.2 71.8 53.1 49.7 25.5
B-16 DataComp-1B 791.4 1400.0 73.5 64.4 66.0 83.6 63.7 60.4 48.4
B-16 OpenAI 784.6 400.0 68.3 58.6 61.9 77.7 55.3 48.2 49.9
B-16 EntityNet 36.0 32.7 66.2 47.4 59.2 64.1 40.9 48.9 23.9
L-14 OpenAI 3328.4 400.0 75.5 71.4 69.9 87.9 69.0 59.6 70.7
L-14 DataComp-1B 3338.6 1400.0 79.2 74.9 72.0 90.8 74.3 68.0 69.6
L-14 DFN-2B 3338.6 2000.0 81.4 74.8 74.6 90.0 74.1 68.3 66.8
H-14 DFN-5B 22164.0 5000.0 83.4 76.3 77.4 93.0 68.4 72.8 69.9
B-32 DataComp-1B (Base model) 295.4 1400.0 69.2 56.3 60.8 78.2 55.2 56.8 30.5
B-32 $\star$ EntityNet 13.3 32.7 69.5 53.3 61.9 74.2 47.9 56.8 25.6
B-32 $\star$ EntityNet (only living organisms) 4.2 10.2 48.2 33.3 43.1 56.7 19.1 32.1 15.8
B-16 DataComp-1B (Base model) 791.4 1400.0 73.5 64.4 66.0 83.6 63.7 60.4 48.4
B-16 $\star$ EntityNet 36.1 32.7 73.5 61.0 66.5 79.0 56.6 59.8 42.9
B-16 $\star$ EntityNet (only living organisms) 11.3 10.2 51.4 39.2 46.1 62.1 25.9 35.1 26.8
B-32 EntityNet Everything 13.1 32.7 61.5 41.0 53.6 58.8 32.6 45.0 14.9
B-32 EntityNet (no living organisms) 9.0 22.5 39.2 28.0 33.6 37.7 29.1 32.0 7.5
B-32 EntityNet (only living organisms) 4.1 10.2 36.0 21.0 31.5 39.6 8.1 17.6 8.1
B-32 EntityNet (no attribute queries) 8.7 21.8 54.8 33.8 47.9 49.2 24.4 36.9 10.5
B-32 EntityNet 50% alt text 13.1 32.7 61.5 41.0 53.6 58.8 32.6 45.0 14.9
B-32 EntityNet 100% alt text 13.1 32.7 59.1 38.1 51.1 55.6 30.1 40.8 13.1
B-32 EntityNet 0% alt text 13.1 32.7 55.7 35.5 48.0 53.0 27.1 36.8 12.4
B-32 EntityNet Full size 13.1 32.7 61.5 41.0 53.6 58.8 32.6 45.0 14.9
B-32 EntityNet 1/2 size 6.6 16.4 54.1 33.6 47.2 51.0 24.4 36.5 8.9
B-32 EntityNet 1/4 size 3.3 8.2 45.2 25.7 39.1 39.9 18.1 25.6 5.9
B-32 EntityNet 1/8 size 1.6 4.1 33.3 17.7 28.5 28.7 12.0 15.3 4.2
B-32 EntityNet 1/16 size 0.8 2.0 19.9 10.0 17.2 17.6 6.8 5.7 2.8
B-32 EntityNet batch size 2,048 13.1 32.7 59.7 40.4 52.5 59.4 31.3 44.5 14.4
B-32 EntityNet batch size 4,096 13.1 32.7 60.9 41.0 53.0 60.0 32.1 44.9 14.9
B-32 EntityNet batch size 8,192 13.1 32.7 61.5 41.0 53.6 58.8 32.6 45.0 14.9
B-32 EntityNet batch size 16,384 13.2 32.7 60.5 39.7 53.2 58.0 30.6 42.9 13.8
B-32 EntityNet batch size 32,768 13.3 32.7 58.6 37.2 51.3 54.6 27.9 40.6 11.6
B Qualitative examples of our dataset
We show randomly sampled images and corresponding textual information of our dataset in Figures 4, 5, 6, and 7. We show an example of our text label sampling strategy in Table 10.
Refer to caption Entity + attribute running polecat A wonderful polecat in its woodland surroundings Polecats Unveiled: Sleek Predators in the Countryside (Mustela Putorius) - Glenlivet Wildlife Polecats Unveiled: Sleek Predators in the Countryside (Mustela Putorius) Black Polecat Photos and Premium High Res Pictures - Getty Images Do Cats Eat Ferrets – What You Should Know! – FAQcats.com Do Cats Eat Ferrets – What You Should Know!
Refer to caption Entity + attribute summer Canada goose Canada Geese Goose Branta - Free photo on Pixabay - Pixabay Canada Geese Goose Branta · Free photo on Pixabay Facts about geese Canada Geese Goose Branta Free Photo On Pixabay Pixabay, 45% OFF
Refer to caption Entity + attribute wild tortoise Greek Tortoise Testudo Graeca Hiding Shell Stock Photo 1425661328 | Shutterstock Elongated Tortoise Indotestudo Elongata Yellow Tortoise Stock Photo 1463951543 | Shutterstock
Refer to caption Entity Orbea decaisneana Orbea decaisneana subs. hesperidum f. cristata
Refer to caption Entity + attribute old walrus What Is A Walrus? What is a Walrus - Walrus Habitat and Behavior - Wild Focus Expeditions Portrait of an old bull walrus resting on his teeth, tooth walker
Figure 4: Randomly sampled images from the EntityNet living subset together with query type, search query, and alt texts (separated by new lines).
Refer to caption Entity Junín red squirrel Curious Eurasian Red Squirrel, Sciurus Vulgaris, Running and Jumping …
Refer to caption Entity chile pine Araucaria araucana - Wikipedia Araucaria araucana - Wikipedia, la enciclopedia libre Araucaria araucana - Wikipedia | Trees to plant, Denver botanic gardens … a tall tree in the middle of a forest Araucaria araucana - Wikipedia | Denver botanic gardens, Outdoor plants … Monkey Puzzle Plant main Monkey Puzzle Plant Care & Growing Basics: Water, Light, Soil, Propagation etc.
Refer to caption Entity Cedronella canariensis Cedronella canariensis
Refer to caption Entity Barbuda Warbler In the face of elite tourism projects, the Barbuda Warbler isn’t the only one that might lose its home Barbuda Warbler - Setophaga subita - Birds of the World / - Barbuda Warbler
Refer to caption Natural type + attribute long neck animal The Strange Elegance of the Giraffe-Necked Antelope | The Ark In Space
Figure 5: Randomly sampled images from the EntityNet living subset together with query type, search query, and alt texts (separated by new lines).
Refer to caption Natural type + attribute carpaccio with lemon Tuna Carpaccio with Fennel and Lemon Recipe - Great British Chefs Tuna carpaccio with fennel and lemon
Refer to caption Natural type + attribute index finger nail A Macro Of Index Finger Nail Stock Footage SBV-301021776 - Storyblocks
Refer to caption Entity akaogiite Akaogiite Cut Out Stock Images
Refer to caption Entity public bookcase manufacta est: public bookcase
Refer to caption Entity BK 117 Eurocopter-Kawasaki BK-117B-2 - DRF - Deutsche Rettungsflugwacht | Aviation Photo #1053819 | Airliners.net Eurocopter-Kawasaki BK-117B-2 - DRF - Deutsche Rettungsflugwacht
Figure 6: Randomly sampled images from the EntityNet non-living subset together with query type, search query, and alt texts (separated by new lines).
Refer to caption Entity cowboy boots Ariat Heritage Rough Stock Cowboy Boots - Square Toe - Country Outfitter
Refer to caption Entity + attribute nacre jewelry Genuine Nacre Necklace Pearl Jewelry Mother Of Pearls Beads | Etsy
Refer to caption Entity + attribute small artificial pond Small backyard pond decoration. Artificial pond in garden. Pool aquatic plants. Pond border decoration. photo Small backyard pond decoration. Artificial pond in garden. Pool aquatic plants. Pond border decoration. Small backyard pond decoration. Artificial pond in garden. Pool aquatic plants. Pond border decoration. 9562060 Stock Photo at Vecteezy Small backyard pond decoration. Artificial pond in garden. Pool aquatic plants. Pond border decoration. Free Photo
Refer to caption Entity + attribute berimbau at rest Premium AI Image | Best Berimbau With Handle Isolated On White Background
Refer to caption Entity + attribute assembled motherboard Red Team rocking: Build the ultimate AMD gaming PC | PCWorld amd fx cpu
Figure 7: Randomly sampled images from the EntityNet non-living subset together with query type, search query, and alt texts (separated by new lines).
Table 10: Example of our text label sampling strategy for an image returned from the entity query of zipper. Probability mass is split 50/50 between image alt texts and texts from the knowledge graph. Between alt texts, we chose uniformly. Between knowledge graph texts we chose the search query 25% of the time, a description 10% of the time (uniformly between descriptions), and an alias otherwise (uniformly between all aliases).
Text Chance Source
Zipper PNG 25% Alt text
yellow zipper PNG image 25% Alt text
zipper 12.5% Search query
zip 5.5% Alias
dingy 5.5% Alias
clasp locker 5.5% Alias
fly 5.5% Alias
zip fastener 5.5% Alias
device for fastening the edges of an opening of fabric or other flexible material 2.5% Description
A device used for fastening, typically made of physical material. 2.5% Description
C Hyperparameter settings
We report hyperparameters for our experiments in Table 11. Similar to Li et al. Li et al. [17], we reduce the context size of the text encoder down from 77 to 32 to reduce VRAM and training time requirements. For a fair comparison with other CLIP models, we report all training cost and training duration as if the training was run at a context length of 77.
D Image Search APIs
Google The Google Image Search API is available via the Google Cloud Platform, and requires an existing programmable search engine to function. It returns up to 10 images per request and page with a limit of 10 pages, i.e., 100 images per query. It costs 5\$ per 1,000 API calls, resulting in costs of about 500\$ to download 1M images. We found the search results from the Google API to be quite different, and arguably worse, from the ones returned when using the regular Google image search. For all our API requests we set the parameter imgColorType to color, imgType to photo, lr to lang_en, and excludeTerms to drawing clipart illustration cartoon vector painting. This way we get mostly real-world images in our search results. We additionally add all aliases and the natural type of the sought entity to the orTerms parameter for entity and entity-attribute queries. Because the Google API returns only up to 10 images per request and page, we search for the following number of pages: 2 pages each for entity queries, 4 pages each for entity-attribute queries, and 10 pages each for all natural-type-attribute queries. We started our search with queries from the living subset on both the Google and Bing APIs. We found the quality and value-for-money ratio of the Bing API to be better, and therefore switched to only using Bing for all other queries.
Bing The Bing Image Search API is available via Microsoft Azure. It returns up to 150 images per request and has no restrictions on the number of accessible pages. It costs 18\$ per 1,000 API calls, resulting in costs of about 120\$ to download 1M images. In our experience, the returned images closely match the ones from the regular Bing image search. For all our API requests we set the parameter imageType to Photo and color to ColorOnly. Unlike the Google API, Bing does not have a way to specify orTerms via a separate request parameter, so we add the natural type of the sought entity to the search query directly, e.g., we search for jaguar animal. The Bing API returns 150 images per request and page, we request one page for each query.
E Querying entities with SPARQL
The SPARQL query over Wikidata used to harvest all entities under a super-entity is displayed in Figure 8. It returns a list of entities from a specified target domain as defined by one or more super-entities. The super-entities can be determined manually by searching for appropriate entities on the Wikidata website. For example, if we want to build a dataset about vehicles, we can set the super-entity to vehicle (Q42889), as done in Table 12. We list the super-entities we considered for our dataset and the relevant statistics in Table 13 for included and Table 14 for excluded super-entities.
Table 11: Hyperparameters used for training and finetuning.
Dataset Model Batch Size Max LR
Weight
Decay
Epochs
Warmup
eps Beta 1 Beta 2
CC12M ViT-B/32 8k 5e-4 0.2 18 2 1e-8 0.9 0.98
CC12M ViT-B/16 8k 5e-4 0.2 18 2 1e-8 0.9 0.98
Ours ViT-B/32 8k 5e-4 0.2 18 2 1e-8 0.9 0.98
Ours ViT-B/16 8k 5e-4 0.2 18 2 1e-8 0.9 0.98
Ours, finetuning ViT-B/32 32k 5e-5 0.2 18 2 1e-8 0.9 0.98
Ours, finetuning ViT-B/16 32k 5e-5 0.2 18 2 1e-8 0.9 0.98
PREFIX wdt: <http://www.wikidata.org/prop/direct/>PREFIX wd: <http://www.wikidata.org/entity/>PREFIX rdfs: <http://www.w3.org/2000/01/rdf-schema#>PREFIX schema: <http://schema.org/>PREFIX wikibase: <http://wikiba.se/ontology#>PREFIX skos: <http://www.w3.org/2004/02/skos/core#>SELECT DISTINCT  ?ent  ?label  ?desc  ?links  (GROUP_CONCAT(DISTINCT ?alias; SEPARATOR=";;;") AS ?aliases)WHERE {  VALUES ?typ { wd:Q42889 }  ?ent wdt:P279* ?typ .  ?ent rdfs:label ?label .  FILTER(LANG(?label) = "en")  ?ent ^schema:about/wikibase:sitelinks ?links .  FILTER(?links >= 5)  OPTIONAL {    ?ent schema:description ?desc .    FILTER(LANG(?desc) = "en")  }  OPTIONAL {    ?ent skos:altLabel ?alias .    FILTER(LANG(?alias) = "en")  }}GROUP BY ?ent ?label ?desc ?linksORDER BY DESC(?links)
Figure 8:  Generic SPARQL query for extracting entities from Wikidata that are related to a given set of super-entities. The super-entities are manually set within the VALUES ?typ {} clause. In this example it is the motor car entity wd:Q42889. A minimum number of sitelinks can also be specified to filter out unpopular entities, here it is set to 5.
Table 12: Vehicle entities and accompanying information as extracted from the Wikidata knowledge graph. Showing the first 5 and last 5 out of 17,015 entities. Note that we only collect entities with sitelinks $\geq 5$. The corresponding SPARQL query is shown in Figure 8.
Identifier Name Description Sitelinks Aliases
Q1420 motor car motorized road vehicle designed to carry one to eight people rather than primarily goods 237 auto / motor vehicle / motor cars / motorcar / cars / car / automobiles / automobile / autocar
Q11442 bicycle pedal-driven two-wheel vehicle 203 bike / Bicycles / cycle / pushbike / pedal cycle / pedal bike
Q197 airplane powered fixed-wing aircraft 196 airplane, aeroplane, plane / powered fixed-wing aircraft / planes / plane / aeroplane / fixed-wing powered aircraft / fixed-wing airplane / aeroplanes / fixed-wing aeroplane / airplanes
Q870 train form of rail transport consisting of a series of connected vehicles 193 rail-train / trains / railway train / railtrain / rail train / railroad train
Q11446 ship large buoyant watercraft 178 marine vessel / vessel / water vessel / ships
Q813876 Bedford JJL motor vehicle 5
Q7077241 Odakyu 20000 series RSE Japanese electric multiple unit trainset 5 RSE / Romancecar RSE / Resort Super Express / Odakyu Romancecar RSE / 20000 series
Q812263 Bavarian Pt 2/3 class of 97 German 2-4-0T locomotives 5 ÖBB 770 / DR Class 70.0 / DRG Class 70.0
Q9177196 Bombardier CRJ1000 regional jet airliner 5 CRJ1000
Q812260 Bavarian PtL 2/2 class of 6+29+13 German 0-4-0T locomotives 5 DB Class 98.3 / DRG Class 98.3 / ÖBB 688
Table 13: The super-entities for building our EntityNet dataset to describe the visual world. The aliases column refers to the set of all aliases collected from the entities. The numbers in this table are slightly higher than the ones we report in the main paper, because they refer to the raw counts of entities and aliases before profanity filtering and the removal of entities that return no results in the image search.
Super-entity Description Examples Entities Aliases
product Anything that can be offered to a market banh mi, navigation system, PlayStation 2 63,676 144,715
substance Any composed matter whose origin is either biological, chemical, or mineral solid lubricant, Chinese tea, eye cups 34,259 111,383
physical tool Physical item that can be used to achieve a goal Patient lift, police transport, instant camera 32,727 71,227
animal Kingdom of multicellular eukaryotic organisms saw-scaled viper, Sporathraupis cyanocephala, Rufous mouse-eared bat 28,000 76,408
plant Living thing in the kingdom of photosynthetic eukaryotes Whitebark Pine, Eucalyptus coccifera, wig knapweed 28,000 55,925
material Substance that can occur in different amounts, all with some similar [mixture of some] characteristics, and with which objects can be made dietary proteins, stone slab tomb, safflower oil 18,021 40,822
vehicle Mobile machine used for transport, whether it has an engine or not, including wheeled and tracked vehicles, air-, water-, and space-craft shipwrecks (objects), Evergreen A-class container ship, VTOL aircraft 17,015 37,849
geographical feature Components of planets that can be geographically located hydrothermal Vents, grooves, street lamp 8,683 19,030
food Any substance consumed to provide nutritional support for the body; form of energy stored in chemical coffee milk, tikka, Friesian Clove 8,464 15,332
architectural structure Human-designed and -made structure rock temples, summerhouse, house of worship 4,507 10,354
anatomical structure Entity with a single connected inherent 3d shape that’s created by coordinated expression of the organism’s own dna bronchi, maxillary wisdom tooth, turtle shell 4,394 9,999
facility Place, equipment, or service to support a specific function public toilet, automobile servicing shop, industrial park 2,767 6,740
physical activity Human physical activity consisting of voluntary bodily movement by skeletal muscles American rules football, archery, water-skiing 2,228 4,422
clothing Covering worn on the body blucher shoe, G-suit, one-piece swimsuit 1,929 4,313
building Structure, typically with a roof and walls, standing more or less permanently in one place shoestore, family restaurant, factory outlet 1,655 3,964
musical instrument Device created or adapted to make musical sounds electroencephalophone, Chinese flutes, oboe 1,450 3,493
organ Collection of tissues with similar functions nasal bone, cranial nerves, ulnar collateral ligament of elbow 1,155 2,450
furniture Movable objects used to equip households, offices, or shops for purposes such as storage, seating, sleeping faldstool, airline seat, bicycle parking rack 388 933
body of water Any significant accumulation of water, generally on a planet’s surface dammed lake, deep-sea hydrothermal vent, marshland 379 792
weather State of the atmosphere cold snap, tropical cyclone, sea of fog 151 304
precipitation Liquid or solid water that falls to the ground hail, thunderstorm, snowfall 43 72
Total Before deduplication 259,891 620,527
Total After deduplication 146,985 368,062
Total After deduplication, without animals and plants 90,985 235,795
Table 14: We consider these super-entities either non-visual, irrelevant, or too specific and do not select related entities when building our dataset.
Super-entity Description
abstract entity entity that does not have a physical existence, including abstract objects and properties
astronomical object physical body of astronomically-significant size, mass, or role, naturally occurring in a universe
city large human settlement
concept semantic unit understood in different ways, e.g. as mental representation, ability or abstract object (philosophy)
continent large landmass identified by convention
country distinct territorial body or political entity
historical event particular incident in history that brings about a historical change
history past events and their tracks or records
imaginary character character known only from narrations (fictional or in a factual manner) without a proof of existence; includes fictional, mythical, legendary or religious characters and similar
language particular system of communication, often named for the region or peoples that use it
language structured system of communication
medical procedure process of medicine done to heal; course of action intended to achieve a result in the delivery of healthcare
organization social entity established to meet needs or pursue goals
planet celestial body directly orbiting a star or stellar remnant
religion social-cultural system
representation entity or process that portrays something else, usually in a simplified or approximated manner
role social role with a set of powers and responsibilities within an organization
science systematic endeavor that builds and organizes knowledge, and the set of knowledge produced by this system
social system patterned series of interrelationships existing between individuals, groups, and institutions
speciality field limited to a specific area of knowledge; specialization in an occupation or branch of learning; a specific use
star astronomical object consisting of a luminous spheroid of plasma held together by its own gravity
temporal entity thing that can be contained within a period of time, or change in state (e.g. events, periods, acts)
work of art aesthetic item or artistic creation; object whose value is its beauty only, not practical usefulness
written work any work expressed in writing, such as inscriptions, manuscripts, documents or maps